### v1

# Over-the-Air Federated Adaptive Data Analysis: Preserving Accuracy via Opportunistic Differential Privacy

Abstract
Adaptive data analysis (ADA) involves a dynamic interaction between an analyst and a dataset owner, where the analyst submits queries sequentially, adapting them based on previous answers. This process can become adversarial, as the analyst may attempt to overfit by targeting non-generalizable patterns in the data. To counteract this, the dataset owner introduces randomization techniques, such as adding noise to the responses. This noise not only helps prevent overfitting but also enhances data privacy. However, it must be carefully calibrated to ensure that the statistical reliability of the responses is not compromised. In this paper, we extend the ADA problem to the context of distributed datasets. Specifically, we consider a scenario where a potentially adversarial analyst interacts with multiple distributed responders through adaptive queries. We assume that the responses are subject to noise introduced by the channel connecting the responders and the analyst. We demonstrate how, through a federated mechanism, this noise can be opportunistically leveraged to enhance the generalizability of ADA, thereby increasing the number of query-response interactions between the analyst and the responders. We illustrate that careful tuning of the transmission amplitude, based on the theoretically achievable bounds, can significantly impact the number of accurately answerable queries.
I Introduction
In the era of big data, the necessity for systems that can dynamically and efficiently analyze vast and complex datasets has never been greater. One prominent approach in this domain is Adaptive Data Analysis (ADA), where the analyst aims to derive insights about an unknown distribution over a given domain by adaptively exchanging a series of query-answer pairs with a responding mechanism that possesses a dataset of samples drawn from that distribution. However, theoretical considerations of ADA in contexts involving distributed datasets is relatively unexplored, despite the growing prevalence of such datasets in modern computing environments. This paper suggests that implementing ADA within distributed data centers can be seen as an extension of Federated Learning, which traditionally focuses on using gradient descent queries to train deep neural networks across decentralized locations. The following subsections offer a necessary examination of ADA and Federated Learning.
I-A Adaptive Data Analysis
In the basic model of ADA, there is one analyst ($\mathcal{A}$) and one answering mechanism ($\mathcal{M}$), communicating with each other through error-free channels with unlimited capacities. Hence, there is no concern about the practical communication limitations in this basic form of the problem. It’s assumed that the answering mechanism $\mathcal{M}$ has access to the dataset $S$ with $n$ samples generated from the domain $\mathcal{X}$ by the distribution $\mathcal{D}_{\mathcal{X}^{n}}$. Albeit, in the general case, $\mathcal{D}_{\mathcal{X}^{n}}$ is not necessarily a product distribution but in the sequel we add this assumption, as it’s the case in most of the basic papers in the field. So we assume that $\mathcal{D}_{\mathcal{X}^{n}}\equiv\mathcal{P}^{n}$, where $\mathcal{P}$ is a distribution on $\mathcal{X}$. The analyst $\mathcal{A}$ doesn’t have direct access to $S$ but can ask queries about it from $\mathcal{M}$. Various families of queries are considered in the ADA papers, among which the most basic one is the statistical queries, studied in the pioneer work [1] and also in several subsequent works such as [2, 3]. In this paper we work with this family of queries, too. The statistical queries are defined in the below.
Definition I.1.
A statistical query is uniquely specified by a function like $q:\mathcal{X}\rightarrow[0,1]$ and its answer is the linear average of the random samples, drawn from a particular distribution. More exact, It’s answer with respect to the distribution $\mathcal{P}$ will be
$q(\mathcal{P})=\mathbb{E}_{x\sim\mathcal{P}}\left[q(x)\right].$ (1)
Despite their simplicity, the statistical queries cover a substantial range of practical queries. For example, gradient descent updates are indeed a vector of statistical queries. Note that the most of federated learning papers take their attention on the gradient descent queries, because of their pivotal role in the learning procedure of deep neural networks. Thus, from another perspective, our work can be seen as an abstraction of those studies.
The analyst $\mathcal{A}$ can make his queries in an adaptive manner based on the previous question-answers. More exact, the $i$’th query, i.e. $q_{i}$, is asked considering the sequence $(q_{1},a_{1},\dots,q_{i-1},a_{i-1})$, in which $a_{j}$ is $\mathcal{M}$’s answer to the $j$’th question, using the dataset $S$. It’s desired to keep the empirical answers close to the accurate answers based on the underlying distribution. The following definition, gives the exact formulation of the considered accuracy criterion.
Definition I.2.
We say that the adaptive data analysis procedure is $(\alpha,\beta)$-accurate if:
$\Pr\left\{\max_{i}{\left|q_{i}(\mathcal{P})-a_{i}\right|}\geq\alpha\right\}\leq\beta,$ (2)
in which $q_{i}(\mathcal{P})$ is the expected value of the answer to the $i$’th query and the probability is calculated on all of the randomnesses of the model.
In the basic model, because $\mathcal{M}$ does not know the action model of $\mathcal{A}$, it considers the worst-case scenario. Specifically, $\mathcal{M}$ assumes that $\mathcal{A}$ selects queries adversarially, aiming to overfit to the dataset $S$ and thereby violate the accuracy criterion defined by (2). On the other hand, $\mathcal{M}$ aims to answer as much query as possible provided that (2) stays valid. So he looks for an efficient answering mechanism to achieve this goal. If the queries were non-adaptive, i.e. chosen independent of observations from data, a natural unbiased action by $\mathcal{M}$ would be to respond with the empirical answers calculated on $S$. But there is well-known examples which shows this can be an inefficient choice for the adaptive situation. For example see [1]. A widely adopted strategy in the research community for handling the adaptive case is to randomize empirical answers by introducing stochastic noise. The concept of adding noise originates from the data privacy framework, particularly differential privacy, which functions as a privacy-enhancing mechanism [4]. In this solution, i.i.d. noise samples from an appropriate distribution are generated and added to each empirically calculated answer. The noise distribution is typically chosen to be either Gaussian or Laplace. In this paper, we proceed with the Gaussian distribution, as we aim to leverage Gaussian channel noise as the source of randomization. The variance of the randomization noise is calibrated to ensure that the accuracy criterion in (2) is satisfied.
Since 2015, numerous papers have explored various aspects of Adaptive Data Analysis (ADA). In this study, we will not delve into the specifics of their theoretical findings; instead, our focus is on applying one of these foundational results to our proposed approach. To the best of our knowledge, no prior work has addressed the physical channel aspects of this issue. Furthermore, this study is the first to investigate the federated scenario within the context of ADA.
I-B Federated learning
As mentioned earlier, previous studies have not examined the physical and distributed aspects of ADA. However, the concept of distributed responders in our work is closely related to federated learning (FL). Accordingly, we reference selected works from the extensive FL literature that address aspects similar to those discussed in our research.
In Federated Learning (FL), a central node, referred to as the Parameter Server (PS), communicates with multiple edge nodes known as clients. All nodes collaborate toward the shared goal of learning the weights of a neural network. Each client has access to a local dataset and contributes to the training process using this data. At the beginning of each training round, the PS broadcasts the latest set of weights to all nodes. Each node then computes the gradient of the loss function based on the current parameter values and its local dataset. Finally, the nodes send their local gradient estimations to the PS, which aggregates them to produce a global, more accurate estimation.
In the literature, various channel models and communication methods have been proposed for the gradient transmission step. These communication approaches are generally categorized into two groups: analog transmission and digital transmission. In this paper, we adopt the analog approach. In analog transmission, clients simultaneously send their estimates to the PS using uncoded transmission. This method is bandwidth-efficient due to the shared use of the channel. Fortunately, due to the additive property of the medium, local estimations naturally aggregate on the air. While this offers computational and storage efficiency, the aggregated signal may carry more noise and interference compared to digital transmission. However, these effects are not always detrimental. Some federated learning studies have leveraged channel distortions to enhance client data privacy. Moreover, [5] shows that although channel imperfections may slow the convergence rate, they can improve the accuracy of the trained network. Thus, there appears to be a dual relationship between data privacy and algorithm accuracy.
Another crucial aspect in FL context is the channel model. One key factor is the Channel State Information (CSI). In [6, 7, 8, 9], perfect global CSI for the PS and perfect local CSI for the edge devices are assumed. In [10, 11, 12, 13, 14, 15], perfect local CSI at the edge devices is assumed. In [16], each edge device is assumed to have imperfect knowledge of its channel gain to the PS, meaning there is some mismatch between its estimation and the exact channel gain. Another type of imperfect CSI is explored in [15, 17], where the PS is assumed to know only the sum of the fading coefficients, and the edge devices have no CSI knowledge. In the scheme proposed by [5], no CSI is required at all.
Another significant factor is the channel’s fading model. The most basic model, adopted by works such as [18, 19], uses the Gaussian MAC model, which does not account for the channel’s attenuating behavior.
For a more practical approach, several works, including [13, 10, 12], consider fading effects, specifically the block-flat fading regime, where the channel condition remains constant during each communication round and may change afterward. In [6], fast fading is considered. Also [11] incorporates both path loss and small-scale fading into the model and assumes that the EP’s have perfect knowledge about both the losses.
I-C Contributions and Paper Structure
In this study, we consider, for the first time, performing ADA through noisy communication channels. As previously mentioned, in the ADA framework, independent noise samples are typically added to answers to maintain statistical validity. Therefore, rather than viewing channel noise as a limitation, we can opportunistically leverage it as a substitute for artificial noise.
Furthermore, we extend this approach to a distributed scenario, where an analyst—potentially adversarial—interacts with multiple EPs via a sequence of queries and responses, rather than working with a single answerer. The responses are transmitted through an Additive White Gaussian Noise (AWGN) channel. We propose a response mechanism that exploits the channel noise to enhance the generalizability of ADA and optimize the number of query-response interactions.
Section II outlines the system model, while Section III presents the core theoretical materials. Section IV examines the ADA problem with a single dataset, considering the effects of the AWGN channel. Section V expands the discussion to ADA involving multiple distributed datasets. Finally, Section VI explores potential future directions and concludes the paper.
II System Model
As depicted in Fig. 1, a distributed situation of ADA is considered. In this problem we assume that each of the $L$ edge points $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\dots$, and $\mathcal{M}_{L}$ are connected to a central point named $\mathcal{A}$ via a communication channel with a limited capacity. Each edge point (EP), e.g. $\mathcal{M}_{l}$, holds a local dataset $S_{l}$ including $n_{l}$ data samples and can act as an answering mechanism. The central point (CP) broadcasts its queries via error-free channels to all or some of them and receives the answers from them through a wireless MAC. Note that, this channel model is similar to the one proposed in lots of recent federated learning works, among which we refer to some of them in the previous section. This scenario can model a network of sensors communicating with a central point, which aim to get some inferences about the environment via data samples collected by the sensors.
More exact, in our considered scheme, the answering EPs transmit an analog signal with amplitude proportional to the answer of query and let the noisy channel to play the role of the stabilizing mechanism. Due to the independence of the noise samples, in each round of the transmission, the answers experience a new randomness, making the situation consistent with the popular solution for controlling the bias in the basic model of adaptive data analysis problem. This idea is also applied in some recent federated learning works to preserve the data privacy of the EPs.
[…]
Figure 1: We have $L$ responders that respond to analyst queries through an AWGN channel.
III Point-to-Point ADA without any communication constraint
In this section we begin with quoting a key theorem about the basic model of the ADA from [3]. This theorem relates the variance of the Gaussian noise to the accuracy parameters of the answering mechanism. Note that our representation of the theorem is a bit different from the reference, but equivalent to it. In the next sections we will apply this theorem in our proposed answering scheme for the federated adaptive data analysis situation.
Theorem III.1 (Theorem 2.1. of [3]).
Fix a Gaussian mechanism parameter $\sigma>0$ and a desired confidence parameter $10<\beta<$ 1. The Gaussian mechanism can be used to answer $k$ statistical queries while satisfying $(\alpha,\beta)$-distributional accuracy, where $\alpha$ is derived from the following expression:
$\alpha=\max\left\{\sqrt{\frac{2}{n\beta}.\min_{\lambda\in\left[0,1\right)}f\left(\lambda\right)}\,,\,\sqrt{8\sigma^{2}\ln{\frac{4k}{\beta}}}\right\},$ (3)
where
$f\left(\lambda\right)\triangleq\frac{1}{\lambda}\left(\frac{k}{n\sigma^{2}}-\ln\left(1-\lambda\right)\right).$ (4)
The first and second term in (3) shows a tradeoff in $\sigma$. The first term which is decreasing in $\sigma$, is due to over-leakage regime, where more variance for noise is required to provide better level of differential privacy. The second term which is decreasing in $\sigma$ represents under-leakage regime, where high variance of noise deteriorates data utility. In the following we derive a more manageable form for (3). First, we resolve the direct functionality of $\alpha$ in $k$, $n$ and $\sigma^{2}$. Then, we manipulate the form of this function to have $k$ as a function of $n$ and $\sigma^{2}$ for a desired given pair of $(\alpha,\beta)$. This latter form, will be directly applicable in our problem setup in the following sections.
Lemma III.1.
Consider the optimization problem $\min_{\left[0,1\right)}{f(\lambda)}$ which is a part of the left expression in (3). The solution of this problem can be expressed in a closed form as follows:
$\frac{W\left(-e^{-(c+1)}\right)\cdot\left(c+\ln\left(-W\left(-e^{-(c+1)}\right)\right)\right)}{1+W\left(-e^{-(c+1)}\right)},$ (5)
where $W\left(\cdot\right)$ is the Lambert function and $c=\frac{k}{n\sigma^{2}}$. We will denote (5) by $g(c)$.
Proof.
A simple inspection shows that the minimum is not occurred in the extreme points. Hence, to find the minimum, we take the derivative of $f(\lambda)$ with respect to $\lambda$ and set it to zero:
$f^{\prime}(\lambda)=-\frac{1}{\lambda^{2}}\left(\frac{k}{n\sigma^{2}}-\ln(1-\lambda)\right)+\frac{1}{\lambda}\cdot\frac{1}{1-\lambda}=0.$
Consequently causing
$\frac{1}{1-\lambda}=\frac{1}{\lambda}\left(\frac{k}{n\sigma^{2}}-\ln(1-\lambda)\right).$
By rearranging the equation and defining $c\triangleq\frac{k}{n\sigma^{2}}$, we have:
$\frac{\lambda}{1-\lambda}+\ln(1-\lambda)=c.$
Defining $u\triangleq\frac{1}{1-\lambda}$ and consequently $\lambda=1-\frac{1}{u}$, we obtain:
$\quad u-\ln(u)=c+1.$
Setting $v\triangleq-u$ and exponentiating both sides of the resulting equation:
$\quad ve^{v}=-e^{-(c+1)}.$
The last equation represents that $v$ is equal to Lambert $W$ function at the point $-e^{-(c+1)}$. Note that Lambert $W$ function is defined as the inverse of $y=xe^{x}$ function. Hence it satisfies: $W(z)e^{W(z)}=z$. Finally, by replacing the equivalent values and a few rearrangement, we get:
$\lambda^{*}=1+\frac{1}{W\left(-e^{-(\frac{k}{n\sigma^{2}}+1)}\right)}.$ (6)
For negative real numbers in $[-1/e,0)$, Lambert function has two branches, denoted by $W_{0}$ and $W_{-1}$. So we need to determine that in (6) which of them should be appeared. The branch $W_{0}$ gives values between $-1$ and 00, while $W_{-1}$ gives values in $(-\infty,-1]$. On the other hand, due to the problem constraint $10\leq\lambda<$ 1. Hence, in (6) only the usage of $W_{-1}$ is valid. Furthermore, note that variation domain of argument of $W$ in (6) lies exactly in $[-1/e,0)$ which is domain of $W_{-1}$. Finally, Substituting $\lambda^{*}$ in $f(\lambda)$, gives (5). ∎
Despite its complicated formulation in (5), $g(c)$ has a benign manner in $c$. Figure 2 depicts that for $c>10$ its growth is highly near the linear.
Figure 2: $g(c)$ versus $c$.
Corollary III.1.
Let $\alpha$, $\beta$ and $n$ to be given. Then $k$ is derived from $\sigma$ as follows:
$k(\sigma;n,\alpha,\beta)=\min\left\{k_{1},k_{2}\right\},$ (7)
where $k_{1}$ and $k_{2}$ are as follows:
$k_{1}(\sigma;n,\alpha,\beta)$ $\triangleq n\sigma^{2}g^{-1}\left(\frac{n\alpha^{2}\beta}{2}\right),$ (8)
$k_{2}(\sigma;\alpha,\beta)$ $\triangleq\frac{\beta}{4}e^{\frac{\alpha^{2}}{8\sigma^{2}}}.$ (9)
Proof.
Letting $\sigma$ vary from 00 to greater values, we see that in (3), before a critical value of $\sigma$, the first term is selected and after that, $\alpha$ is determined by the second term. When $\alpha$ is fixed, functionality of $k$ on $\sigma$ can be derived from one of these two terms. For each $\sigma$, the term that gives lower $k$ is the correct term. ∎
In the following sections we set $\alpha=0.1$ and $\beta=0.05$, which are typical values in the literature. For this values, figure 2 shows that when $n>4\times 10^{4}$, the first term is highly near
$\hat{k_{1}}(\sigma;n,\alpha,\beta)=w\cdot\frac{n^{2}\sigma^{2}\alpha^{2}\beta}{2}+bn\sigma^{2},$ (10)
where $w$ and $b$ are some constants.
IV Point-to-Point ADA in the Presence of a Gaussian Channel
At the first step, to make our idea clear, we start with the case in which there is only one EP, answering the PS’s adaptive queries via a Gaussian channel. More exact, at the $i$’th question-answering round of this interactive procedure, PS sends $q_{i}(\cdot)$ through an error-free channel and receives $a_{i}$ as the answer from EP through a Gaussian channel with noise variance of $\sigma^{2}$. Conventionally, noise is considered as a limiting factor in the communication procedures. But in this work we opportunistically employ it as a constructive element. As noted in I-A we aim to exploit the intrinsic channel noise as the randomization noise samples in the Gaussian answering mechanism. In this scheme, at each round, such as round $i$, the EP computes the empirical average of the query as
$q_{i}(S)=\frac{1}{n}\sum_{j=1}^{n}{q_{i}(x_{j})},$ (11)
then multiplies it by $A_{\mathsf{t}}$. The resulting signal is broadcast without additional coding, using analog Pulse Amplitude Modulation (PAM) with a maximum amplitude of $A_{\mathsf{t}}$. Using a matched filter, PS detects a noisy version of the transmitted signal. Without any hard decision, he considers the detected signal as $a_{i}$, i.e. the answer of his asked query. So we have:
$a_{i}=A_{\mathsf{t}}q_{i}(S)+z_{i},$ (12)
where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the sampled Gaussian channel noise at the receiver. Assuming that the channel condition remains unchanged during the whole ADA procedure, it’s clear that $n_{i}$’s are independent and identically distributed. So, the proposed scheme completely resembles the Gaussian answering mechanism in ADA and we can apply theorem III.1 to this case. Actually we use this theorem to set the tunable communication parameters in a way that the desired accuracy level is met. We assume that $(\alpha,\beta)$ has been given and the dataset size, $n$, is fixed. Based on this parameters and as a function of $A_{\mathsf{t}}$, we can derive the maximum number of allowable queries, for which still (2) remains valid.
In the basic model of ADA, it’s assumed that the queries maximum value is fixed and equal to 1. There, the variance of the artificial randomization noise is calibrated in a way that the best performance is achieved. On the other hand, in our proposed model, the noise variance is fixed and determined by the physical later conditions. Hence, here we instead tune the queries maximum amplitude, i.e. $A_{\mathsf{t}}$, for the best performance achievement. Also note that $\alpha$ in (2) is considered under the assumption that $|q_{i}(.)|\leq 1$. Nonetheless, we keep using (2) by normalizing the answers maximum values to 1, when investigating this relation. More exact, the validity condition of (2) for the situation $(\alpha=\alpha_{0}\cdot A_{0},\beta=\beta_{0},A_{\mathsf{t}}=A_{0},\sigma=\sigma_{0})$ is completely equivalent with the situation $(\alpha=\alpha_{0},\beta=\beta_{0},A_{\mathsf{t}}=1,\sigma=\sigma_{0}/A_{0})$. So, by tuning $A_{\mathsf{t}}$ we implicitly calibrate the noise variance in the equivalent normalized model.
Therefore, to apply the theorem III.1, it’s sufficient to replace $\sigma$ in (7), with $\frac{\sigma}{A_{\mathsf{t}}}$. In the figure 3 , $k$ is plotted in $\frac{\sigma}{A_{\mathsf{t}}}$, for some different values of $n$ and given desired accuracy parameters $\alpha=0.1$ and $\beta=0.05$.
Figure 3: Achievable number of accurately answerable queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
Also, in figure 4 the maximum achievable number of queries is drawn versus $n$. In this figure, for each $n$, the best $\sigma/A_{\mathsf{t}}$ that maximizes $k$ is considered. As depicted in the figure, $k_{m}ax$ follows a quadratic growth in $n$. This observation confirms the well known fact in the ADA problem that for the worst case setup, till $O(n^{2})$ number of queries can be $(\alpha,\beta)$-accurately be answered [20].
Figure 4: Maximum achievable number of accurately answerable queries versus $n$ for $\alpha=0.1$ and $\beta=0.05$.
Furthermore, figure 3 shows that as $n$ increases, the optimum value of $\sigma/A_{\mathsf{t}}$ very slowly shifts to the left. This optimum point is the intersection of the quadratically increasing left curve and the exponentially decreasing right curve, , see $k_{1}$ and $k_{2}$ in (8). For our considered dataset sizes, optimum $\sigma/A_{\mathsf{t}}$ approximately lies in $[0.008,0.01]$. Furthermore, when $\sigma/A_{\mathsf{t}}$ tends to zero, the number of accurate queries reduce to values close to zero, too. In this low-noise regime the amount of randomization is not sufficient for ensuring that the dataset will not leak. Consequently, in this case by asking only few queries the analyst can overfit to the dataset and violate the accuracy criterion (2). Hence, in contrast to ordinary communication scenarios, increasing the power transmission may cause non-efficiency in this setup. On the other hand, computation shows that in order to have a non-zero number of $(\alpha,\beta)$-accurate query-answers by this scheme, $\sigma/A_{\mathsf{t}}$ should be lower than 0.017 or equivalently $A_{\mathsf{t}}$ should be approximately 60 times greater than the standard deviation of the noise. This ratio for the optimum situation lies in $[100,125]$. Equivalently, a minimum SNR of 32 dB is required and optimum SNR is in the range [37, 38.9] dB. However, in high-noise regime, considering practical limitations on the transmission power, improving $k$ may not be straight-forward. We show in the next section that federation of some edge answering points can remedy this situation.
V Distributed ADA in the Presence of a Gaussian Channel
In the previous section we considered the point-to-point scenario to make the idea of usefully exploiting the channel noise clear. We showed that by calibrating the amplitude of transmitted signal, achievable number of accurately answered queries can be optimized. Furthermore, we see that optimum $A_{\mathsf{t}}$ depends on the dataset size, $n$. On the other hand, although we’ve succeeded theoretically to utilize the noise as a part of adaptive query answering procedure, but the required SNR range may not be manageable in some of communication situations. This concern becomes more crucial when the answerer is a personal device or a low-power sensor at the edge of network with a long distance from the receiver. As we mentioned previously, this can be a reasonable use case of our proposed setup. In addition, we observed that for a fixed $A_{\mathsf{t}}$ the required dataset size, i.e. $n$, grows by $O(\sqrt{k})$. So, to guarantee more accurate (or private) rounds of data analysis, the edge point needs to answer the queries based on larger datasets. However, it may be impractical for him to gather or store large samples. In this section we show that both of the power and data limitations can be handled in the distributed scenario.
We consider a situation that $L$ Edge Points simultaneously transmit their answers to adaptive queries of the Central point through an analog Gaussian multiple access channel (MAC). For simplicity, let’s assume that the dataset sizes of EPs are the same, i.e. $n_{l}=n_{0}$ for any $l\in\{1,2,\dots,L\}$. Similar to the previous section, each EP, like $\mathcal{M}_{l}$, computes $q_{i}(S_{l})=\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}$ and transmit $A_{\mathsf{t}}q_{i}(S_{l})$ without any coding. Because of the additive nature of the air interface, the receiver gets a noisy version of the aggregated signals. So, the received signal at the CP will be $a_{i}=A_{\mathsf{t}}\sum_{l=1}^{L}{q_{i}(S_{l})}+z_{i}$, in which $z_{i}\sim\mathcal{N}(0,\sigma^{2})$. In order to make an equivalence with the point-to-point case, we consider $a_{i}/L$ in the following. By substituting $q_{i}(S_{l})$, we will have:
$\frac{a_{i}}{L}=\frac{A_{\mathsf{t}}}{Ln_{0}}\sum_{l=1}^{L}{\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}}+\frac{z_{i}}{L}.$ (13)
The above equation completely resembles a point-to-point scenario with $n_{\text{eq}}=Ln_{0}$, $\sigma_{\text{eq}}=\frac{\sigma}{L}$ and $A_{T,eq}=A_{\mathsf{t}}$. We see that by this cooperative transmission scheme, both the equivalent dataset and the noise variance is improved by a factor of $L$. By considering these equivalent parameters, the previous section results can be applied in this scenario, too. Therefore, we can use (7) with the aforesaid $n_{\text{eq}}$ and $\sigma_{\text{eq}}$. In this case, we fix $\sigma/A_{\mathsf{t}}=0.5$ and derive $k$ as a function of $L$, for different values of $n$. Considering the previous section, we know that $L$ should be large enough to make $\sigma_{\text{eq}}$ less than $0.017$ and consequently, attain a positive value for $k$. Figure 5 shows the results for this scenario. The parameters $\alpha$ and $\beta$ are the same with the previous scenario.
Figure 5: Maximum number of accurately answered queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
We see that for each $n$, before a critical number of EPs, $k$ grows exponentially in $L$, which is due to under-leakage regime and is related to $k_{2}$ in (8). This critical $L$ shifts very slowly in $n$ to the right. For each $n$, after this $L$, growth of $k$ becomes very slow such that in figure 5 it seems constant. The reason is that after the critical $L$, $k$ obeys from $k_{1}$ function in (8). As we noted previously, $k_{1}$ can be approximated by $\hat{k_{1}}$ in (10). In this function, the quadratic term is strongly dominant for large $n$. In this quadratic term, $n^{2}\sigma^{2}$ is determining. On the other hand, in the distributed scenario we have
$n_{\text{eq}}\sigma_{\text{eq}}=(Ln)(\sigma/L)=\sigma,$
which does not depend on $L$. This argument also provides a perspective for addressing this inefficient growth stoppage, which is discussed in the following subsection.
V-A Improving number of queries in EP counts
Considering our descriptions on the figure 3 about the optimum value of $\sigma/A_{\mathsf{t}}$, EPs in the distributed scenario need to decrease $A_{\mathsf{t}}$ with respect to $L$, after that $L$ passes the critical value. For large $n_{\text{eq}}$, optimum $\sigma_{\text{eq}}/A_{\mathsf{t}}$ is about $0.008$. The exact value can be derived by setting $k_{1}=k_{2}$ in (7). Naming this value by $s_{\text{opt}}(n_{\text{eq}})$, optimum $\sigma/A_{\mathsf{t}}$ will be $Ls_{\text{opt}}(n_{\text{eq}})$. In other words, EPs should tune their $A_{\mathsf{t}}$ equal to $\frac{\sigma}{Ls_{\text{opt}}(Ln)}$. As we noted, in the previous section, $s_{\text{opt}}$ changes very slowly. Hence, we can approximately say that in optimum regime $A_{\mathsf{t}}\propto 1/L$. Figure 6 depicts the resulted improvement in the growth of $k$, after optimizing $A_{\mathsf{t}}$.
Figure 6: Maximum number of accurately answered queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
VI Future Work and Conclusion
In this work, after making a previous result in ADA more proper for our analyses, we showed that Gaussian channel noise can be used as an opportunity when the basic model of ADA is considered. Of course, in this situation minimum required SNR was considerable. Then, we showed that the federation of the edge answering points can improve the equivalent noise variance and total dataset size by a linear factor when the shared medium is a Gaussian MAC.
One important extension for further study can be considering the fading channel, which is a more realistic model. In this situation, each EP’s transmission experiences a multiplicative distortion before receiving by the CP. In the other word, the received signal during the i’th answering round will be as follows
$Y_{i}=\sum_{l=1}^{L}{h_{il}X_{il}+z_{i}},$ (14)
in which $X_{il}$ is the transmitted signal by the $l$’th EP and $h_{il}$ is the fading coefficient of the wireless channel from this EP to the CP. Based on various scenarios of Channel State Information (CSI) knowledge, different scenarios can be studied: CSI known only at the transmitter, CSI known only at the receiver, and CSI not known at either. Overally, a main challenge in the fading scenario is that the unweighted aggregation of transmissions is not intrinsically prepared anymore.

### v2

@@
 To counteract this, the dataset owner introduces randomization techniques, such as adding noise to the responses.
-This noise not only helps prevent overfitting but also enhances data privacy.
+This noise not only helps prevent overfitting, but also enhances data privacy.
 However, it must be carefully calibrated to ensure that the statistical reliability of the responses is not compromised.
@@
 Specifically, we consider a scenario where a potentially adversarial analyst interacts with multiple distributed responders through adaptive queries.
-We assume that the responses are subject to noise introduced by the channel connecting the responders and the analyst.
-We demonstrate how, through a federated mechanism, this noise can be opportunistically leveraged to enhance the generalizability of ADA, thereby increasing the number of query-response interactions between the analyst and the responders.
-We illustrate that careful tuning of the transmission amplitude, based on the theoretically achievable bounds, can significantly impact the number of accurately answerable queries.
+We assume the responses are subject to noise, introduced by the channel connecting the responders and the analyst.
+We demonstrate how this noise can be opportunistically leveraged through a federated mechanism to enhance the generalizability of ADA, thereby increasing the number of query-response interactions between the analyst and the responders.
+We illustrate that the careful tuning of the transmission amplitude based on the theoretically achievable bounds can significantly impact the number of accurately answerable queries.
 I Introduction
-In the era of big data, the necessity for systems that can dynamically and efficiently analyze vast and complex datasets has never been greater.
-One prominent approach in this domain is Adaptive Data Analysis (ADA), where the analyst aims to derive insights about an unknown distribution over a given domain by adaptively exchanging a series of query-answer pairs with a responding mechanism that possesses a dataset of samples drawn from that distribution.
-However, theoretical considerations of ADA in contexts involving distributed datasets is relatively unexplored, despite the growing prevalence of such datasets in modern computing environments.
-This paper suggests that implementing ADA within distributed data centers can be seen as an extension of Federated Learning, which traditionally focuses on using gradient descent queries to train deep neural networks across decentralized locations.
-The following subsections offer a necessary examination of ADA and Federated Learning.
+In the era of big data, the necessity for systems that can dynamically and efficiently analyze vast and complex datasets has never been greater.One prominent approach in this domain is Adaptive Data Analysis (ADA).
+In ADA, the analyst aims to derive insights about an unknown distribution over a given domain by adaptively exchanging query-answer pairs with a responding mechanism that possesses a dataset of samples drawn from that distribution.
+However, theoretical considerations of ADA in contexts involving distributed datasets are relatively unexplored despite the growing prevalence of such datasets in modern computing environments.This paper proposes implementing ADA in distributed systems as an extension of Federated Learning (FL), which traditionally focuses on using gradient descent queries to train deep neural networks across decentralized locations.
+The following subsections provide a necessary examination of ADA and FL.
 I-A Adaptive Data Analysis
 In the basic model of ADA, there is one analyst ($\mathcal{A}$) and one answering mechanism ($\mathcal{M}$), communicating with each other through error-free channels with unlimited capacities.
-Hence, there is no concern about the practical communication limitations in this basic form of the problem.
-It’s assumed that the answering mechanism $\mathcal{M}$ has access to the dataset $S$ with $n$ samples generated from the domain $\mathcal{X}$ by the distribution $\mathcal{D}_{\mathcal{X}^{n}}$.
-Albeit, in the general case, $\mathcal{D}_{\mathcal{X}^{n}}$ is not necessarily a product distribution but in the sequel we add this assumption, as it’s the case in most of the basic papers in the field.
-So we assume that $\mathcal{D}_{\mathcal{X}^{n}}\equiv\mathcal{P}^{n}$, where $\mathcal{P}$ is a distribution on $\mathcal{X}$.
-The analyst $\mathcal{A}$ doesn’t have direct access to $S$ but can ask queries about it from $\mathcal{M}$.
-Various families of queries are considered in the ADA papers, among which the most basic one is the statistical queries, studied in the pioneer work [1] and also in several subsequent works such as [2, 3].
-In this paper we work with this family of queries, too.
-The statistical queries are defined in the below.
+Hence, there is no concern about practical communication limitations in this basic form of the problem.
+It has assumed that the answering mechanism $\mathcal{M}$ has access to the dataset $S$ with $n$ samples generated from the domain $\mathcal{X}$ by the distribution $\mathcal{D}_{\mathcal{X}^{n}}$.
+Albeit, in the general case, $\mathcal{D}_{\mathcal{X}^{n}}$ is not necessarily a product distribution, in the sequel, we add this assumption as it is the case in most of the basic papers in the field.
+So, we assume that $\mathcal{D}_{\mathcal{X}^{n}}\equiv\mathcal{P}^{n}$, where $\mathcal{P}$ is a distribution on $\mathcal{X}$.
+The analyst $\mathcal{A}$ does not have direct access to $S$, but can query $\mathcal{M}$ about it.Various families of queries are considered in the ADA papers, among which the most basic one is the statistical queries, studied in the pioneering work [1] and also in several subsequent works such as [2, 3].
+In this paper, we also focus on this class of queries, which are formally defined below.
 Definition I.1.
@@
 $q(\mathcal{P})=\mathbb{E}_{x\sim\mathcal{P}}\left[q(x)\right].$ (1)
-Despite their simplicity, the statistical queries cover a substantial range of practical queries.
-For example, gradient descent updates are indeed a vector of statistical queries.
-Note that the most of federated learning papers take their attention on the gradient descent queries, because of their pivotal role in the learning procedure of deep neural networks.
-Thus, from another perspective, our work can be seen as an abstraction of those studies.
-The analyst $\mathcal{A}$ can make his queries in an adaptive manner based on the previous question-answers.
-More exact, the $i$’th query, i.e.
-$q_{i}$, is asked considering the sequence $(q_{1},a_{1},\dots,q_{i-1},a_{i-1})$, in which $a_{j}$ is $\mathcal{M}$’s answer to the $j$’th question, using the dataset $S$.
-It’s desired to keep the empirical answers close to the accurate answers based on the underlying distribution.
-The following definition, gives the exact formulation of the considered accuracy criterion.
+Despite their simplicity, statistical queries encompass a substantial range of practical applications.
+For instance, gradient descent updates can be viewed as a vector of statistical queries.
+Notably, most FL papers focus on gradient descent queries due to their pivotal role in the learning process of deep neural networks.
+Consequently, our work can be regarded as an abstraction of these studies.
+The analyst $\mathcal{A}$ can adaptively make queries based on the previous question-answer pairs.
+More precisely, the $i$th query, $q_{i}$, is formulated considering the sequence $(q_{1},a_{1},\dots,q_{i-1},a_{i-1})$, where $a_{j}$ is $\mathcal{M}$’s answer to the $j$th question, using the dataset $S$.
+The goal is to ensure that the empirical answers remain close to the true answers derived from the underlying distribution.
+The following definition provides the exact formulation of the accuracy criterion.
 Definition I.2.
@@
 In the basic model, because $\mathcal{M}$ does not know the action model of $\mathcal{A}$, it considers the worst-case scenario.
-Specifically, $\mathcal{M}$ assumes that $\mathcal{A}$ selects queries adversarially, aiming to overfit to the dataset $S$ and thereby violate the accuracy criterion defined by (2).
-On the other hand, $\mathcal{M}$ aims to answer as much query as possible provided that (2) stays valid.
-So he looks for an efficient answering mechanism to achieve this goal.
-If the queries were non-adaptive, i.e.
-chosen independent of observations from data, a natural unbiased action by $\mathcal{M}$ would be to respond with the empirical answers calculated on $S$.
-But there is well-known examples which shows this can be an inefficient choice for the adaptive situation.
-For example see [1].
-A widely adopted strategy in the research community for handling the adaptive case is to randomize empirical answers by introducing stochastic noise.
-The concept of adding noise originates from the data privacy framework, particularly differential privacy, which functions as a privacy-enhancing mechanism [4].
-In this solution, i.i.d.
-noise samples from an appropriate distribution are generated and added to each empirically calculated answer.
+Expressly, $\mathcal{M}$ assumes that $\mathcal{A}$ selects queries adversarially, aiming to overfit to the dataset $S$ and thereby violate the accuracy criterion defined by (2).
+On the other hand, $\mathcal{M}$ aims to answer as many queries as possible while ensuring that (2) remains satisfied.
+To achieve this, $\mathcal{M}$ seeks an efficient answering mechanism.
+In the case of non-adaptive queries (i.e., queries selected independently of observations from the data), a natural unbiased approach for $\mathcal{M}$ would be to respond using the empirical answers computed from $S$.
+Nevertheless, there are well-known examples demonstrating that this approach can be inefficient in adaptive scenarios.
+For example, see [1].
+A widely adopted strategy in the research community for addressing the adaptive case is to introduce randomness into empirical answers by adding stochastic noise.
+The idea of adding noise comes from the data privacy framework, particularly differential privacy, which serves as a privacy-enhancing mechanism [4].
+In this approach, i.i.d.
+noise samples from a suitable distribution are generated and added to each empirically computed answer.
+The variance of the noise is calibrated to ensure that the accuracy criterion in (2) is satisfied.
 The noise distribution is typically chosen to be either Gaussian or Laplace.
-In this paper, we proceed with the Gaussian distribution, as we aim to leverage Gaussian channel noise as the source of randomization.
-The variance of the randomization noise is calibrated to ensure that the accuracy criterion in (2) is satisfied.
-Since 2015, numerous papers have explored various aspects of Adaptive Data Analysis (ADA).
-In this study, we will not delve into the specifics of their theoretical findings; instead, our focus is on applying one of these foundational results to our proposed approach.
-To the best of our knowledge, no prior work has addressed the physical channel aspects of this issue.
-Furthermore, this study is the first to investigate the federated scenario within the context of ADA.
+In this paper, we focus on the Gaussian distribution, as we aim to leverage Gaussian channel noise as the source of randomization.
+Since 2015, numerous studies have examined various aspects of ADA.
+In this paper, we apply one of the foundational results to our approach, rather than delving into the theoretical details.
+To our knowledge, no previous work has investigated the physical channel aspects of this problem.
+Moreover, this study is the first to explore the federated scenario within the context of ADA.
 I-B Federated learning
-As mentioned earlier, previous studies have not examined the physical and distributed aspects of ADA.
-However, the concept of distributed responders in our work is closely related to federated learning (FL).
-Accordingly, we reference selected works from the extensive FL literature that address aspects similar to those discussed in our research.
-In Federated Learning (FL), a central node, referred to as the Parameter Server (PS), communicates with multiple edge nodes known as clients.
-All nodes collaborate toward the shared goal of learning the weights of a neural network.
-Each client has access to a local dataset and contributes to the training process using this data.
-At the beginning of each training round, the PS broadcasts the latest set of weights to all nodes.
-Each node then computes the gradient of the loss function based on the current parameter values and its local dataset.
-Finally, the nodes send their local gradient estimations to the PS, which aggregates them to produce a global, more accurate estimation.
-In the literature, various channel models and communication methods have been proposed for the gradient transmission step.
-These communication approaches are generally categorized into two groups: analog transmission and digital transmission.
+As previously noted, prior studies have not explored the physical and distributed aspects of ADA.
+However, the notion of distributed responders introduced in this work closely aligns with concepts in FL.
+To provide context, we cite selected works from the extensive FL literature that explore ideas analogous to those studied here.
+In FL, a central node, called the Parameter Server (PS), interacts with multiple edge nodes, or clients.
+Together, these nodes work to train a shared neural network.
+Each client uses its local dataset to contribute to the training process.
+At the start of each training round, the PS broadcasts the updated weights to all nodes.
+Each node computes the gradient of the loss function using the current parameter values and its local dataset.
+Finally, the nodes send their local gradient estimations to the PS, which aggregates them to produce a more accurate global estimation.
+The literature proposes various channel models and communication methods for the gradient transmission step.
+These approaches are generally divided into two categories: analog transmission and digital transmission.
 In this paper, we adopt the analog approach.
-In analog transmission, clients simultaneously send their estimates to the PS using uncoded transmission.
-This method is bandwidth-efficient due to the shared use of the channel.
-Fortunately, due to the additive property of the medium, local estimations naturally aggregate on the air.
-While this offers computational and storage efficiency, the aggregated signal may carry more noise and interference compared to digital transmission.
-However, these effects are not always detrimental.
-Some federated learning studies have leveraged channel distortions to enhance client data privacy.
-Moreover, [5] shows that although channel imperfections may slow the convergence rate, they can improve the accuracy of the trained network.
-Thus, there appears to be a dual relationship between data privacy and algorithm accuracy.
-Another crucial aspect in FL context is the channel model.
-One key factor is the Channel State Information (CSI).
-In [6, 7, 8, 9], perfect global CSI for the PS and perfect local CSI for the edge devices are assumed.
-In [10, 11, 12, 13, 14, 15], perfect local CSI at the edge devices is assumed.
-In [16], each edge device is assumed to have imperfect knowledge of its channel gain to the PS, meaning there is some mismatch between its estimation and the exact channel gain.
-Another type of imperfect CSI is explored in [15, 17], where the PS is assumed to know only the sum of the fading coefficients, and the edge devices have no CSI knowledge.
-In the scheme proposed by [5], no CSI is required at all.
+With analog transmission, clients simultaneously send their estimates to the PS using uncoded transmission.
+This method is bandwidth-efficient because the channel is shared.
+Moreover, the additive property of the medium allows local estimations to naturally aggregate in the air.
+While this approach offers computational and storage efficiency, the aggregated signal may contain more noise and interference than digital transmission.
+However, these effects are not inherently detrimental.
+Some FL studies have utilized channel distortions to enhance client data privacy.
+Furthermore, [5] demonstrates that, although channel imperfections can slow the convergence rate, they may also improve the accuracy of the trained network.
+This suggests a nuanced relationship between data privacy and algorithm accuracy.
+In the context of FL, the channel model is a critical consideration, particularly the availability of Channel State Information (CSI).
+Several studies, such as [6, 7, 8, 9], assume perfect global CSI at the PS and perfect local CSI at the edge devices.
+Some Others, including [10, 11, 12, 13, 14, 15], focus solely on perfect local CSI at the edge devices.
+In contrast, [16] assumes that each edge device has imperfect knowledge of its channel gain to the PS, resulting in some mismatch between the estimated and actual channel gain.
+Another form of imperfect CSI is examined in [15, 17], where the PS is aware only of the sum of the fading coefficients, and the edge devices lack any CSI knowledge.
+Finally, [5] proposes a scheme that requires no CSI at all.
 Another significant factor is the channel’s fading model.
@@
 In [6], fast fading is considered.
-Also [11] incorporates both path loss and small-scale fading into the model and assumes that the EP’s have perfect knowledge about both the losses.
+Also [11] incorporates both path loss and small-scale fading into the model, assuming that the EPs have perfect knowledge of both types of losses.
 I-C Contributions and Paper Structure
-In this study, we consider, for the first time, performing ADA through noisy communication channels.
-As previously mentioned, in the ADA framework, independent noise samples are typically added to answers to maintain statistical validity.
-Therefore, rather than viewing channel noise as a limitation, we can opportunistically leverage it as a substitute for artificial noise.
-Furthermore, we extend this approach to a distributed scenario, where an analyst—potentially adversarial—interacts with multiple EPs via a sequence of queries and responses, rather than working with a single answerer.
+In this study, we explore, for the first time, the execution of ADA over noisy communication channels.
+As noted earlier, the ADA framework typically introduces independent randomnesses to answers to preserve statistical validity.
+Thus, instead of treating channel noise as a drawback, we propose leveraging it as a natural substitute for artificially added noise.
+Furthermore, we extend this approach to a distributed scenario, where an analyst—potentially adversarial—interacts with multiple EPs via a sequence of queries and responses rather than working with a single answerer.
 The responses are transmitted through an Additive White Gaussian Noise (AWGN) channel.
@@
 II System Model
-As depicted in Fig.
-1, a distributed situation of ADA is considered.
-In this problem we assume that each of the $L$ edge points $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\dots$, and $\mathcal{M}_{L}$ are connected to a central point named $\mathcal{A}$ via a communication channel with a limited capacity.
-Each edge point (EP), e.g.
-$\mathcal{M}_{l}$, holds a local dataset $S_{l}$ including $n_{l}$ data samples and can act as an answering mechanism.
-The central point (CP) broadcasts its queries via error-free channels to all or some of them and receives the answers from them through a wireless MAC.
-Note that, this channel model is similar to the one proposed in lots of recent federated learning works, among which we refer to some of them in the previous section.
-This scenario can model a network of sensors communicating with a central point, which aim to get some inferences about the environment via data samples collected by the sensors.
-More exact, in our considered scheme, the answering EPs transmit an analog signal with amplitude proportional to the answer of query and let the noisy channel to play the role of the stabilizing mechanism.
-Due to the independence of the noise samples, in each round of the transmission, the answers experience a new randomness, making the situation consistent with the popular solution for controlling the bias in the basic model of adaptive data analysis problem.
-This idea is also applied in some recent federated learning works to preserve the data privacy of the EPs.
+As depicted in Figure 1, a distributed situation of ADA is considered.
+In this problem, we assume that each of the $L$ Edge Points (EPs) $\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{L}$ are connected to a Central Point (CP), referred to as the analyst, via a communication channel with limited capacity.
+Each EP, such as $\mathcal{M}_{l}$, holds a local dataset $S_{l}$ containing $n_{l}$ data samples and functions as an answering mechanism.
+The CP broadcasts its queries via error-free channels to all or some of them, and receives the answers from them through a wireless MAC.
+Note that this channel model is similar to those proposed in many recent FL studies, some of which were referenced in the previous section.
+This scenario can model a network of sensors communicating with a CP to infer environmental characteristics using data samples collected by the sensors.
+In our proposed scheme, the answering EPs transmit analog signals with amplitudes proportional to the query answers, allowing the noisy channel to act as a stabilizing mechanism.
+Due to the independence of the noise realizations, each transmission experience a new randomness, aligning with the widely adopted solution for controlling bias in the basic model of the ADA problem.
+This approach has also been applied in some recent FL studies to preserve data privacy for the EPs.
 […]
@@
 III Point-to-Point ADA without any communication constraint
-In this section we begin with quoting a key theorem about the basic model of the ADA from [3].
-This theorem relates the variance of the Gaussian noise to the accuracy parameters of the answering mechanism.
-Note that our representation of the theorem is a bit different from the reference, but equivalent to it.
-In the next sections we will apply this theorem in our proposed answering scheme for the federated adaptive data analysis situation.
+In this section, we present a key theorem from [3] regarding the basic model of ADA.
+The theorem establishes a relationship between the variance of Gaussian noise and the accuracy parameters of the answering mechanism.
+While our representation differs slightly from the original, it remains mathematically equivalent.
+In the subsequent sections, we apply this theorem to our proposed answering scheme for the federated adaptive data analysis.
 Theorem III.1 (Theorem 2.1. of [3]).
@@
 $f\left(\lambda\right)\triangleq\frac{1}{\lambda}\left(\frac{k}{n\sigma^{2}}-\ln\left(1-\lambda\right)\right).$ (4)
-The first and second term in (3) shows a tradeoff in $\sigma$.
-The first term which is decreasing in $\sigma$, is due to over-leakage regime, where more variance for noise is required to provide better level of differential privacy.
-The second term which is decreasing in $\sigma$ represents under-leakage regime, where high variance of noise deteriorates data utility.
-In the following we derive a more manageable form for (3).
-First, we resolve the direct functionality of $\alpha$ in $k$, $n$ and $\sigma^{2}$.
-Then, we manipulate the form of this function to have $k$ as a function of $n$ and $\sigma^{2}$ for a desired given pair of $(\alpha,\beta)$.
+The first and second terms in (3) illustrate a tradeoff in $\sigma$.
+The first term, which decreases with $\sigma$, is due to over-leakage regime, corresponds to the over-leakage regime, where a higher variance for noise is required to achieve a better level of differential privacy.
+The second term, which increases with $\sigma$, represents the under-leakage regime, where a high variance of noise deteriorates data utility.
+In the following, we derive a more manageable form for (3).
+First, we express the direct dependence of $\alpha$ on $k$, $n$ and $\sigma^{2}$.
+Then, we manipulate this function to express $k$ as a function of $n$ and $\sigma^{2}$ for a given pair of $(\alpha,\beta)$.
 This latter form, will be directly applicable in our problem setup in the following sections.
@@
 $f^{\prime}(\lambda)=-\frac{1}{\lambda^{2}}\left(\frac{k}{n\sigma^{2}}-\ln(1-\lambda)\right)+\frac{1}{\lambda}\cdot\frac{1}{1-\lambda}=0.$
-Consequently causing
+Consequently, this results in
 $\frac{1}{1-\lambda}=\frac{1}{\lambda}\left(\frac{k}{n\sigma^{2}}-\ln(1-\lambda)\right).$
@@
 $\quad ve^{v}=-e^{-(c+1)}.$
-The last equation represents that $v$ is equal to Lambert $W$ function at the point $-e^{-(c+1)}$.
-Note that Lambert $W$ function is defined as the inverse of $y=xe^{x}$ function.
-Hence it satisfies: $W(z)e^{W(z)}=z$.
-Finally, by replacing the equivalent values and a few rearrangement, we get:
+The last equation shows that $v$ equals Lambert $W$ function evaluated at the point $-e^{-(c+1)}$.
+The Lambert $W$ function is defined as the inverse of $y=xe^{x}$ function, meaning it satisfies $W(z)e^{W(z)}=z$.
+Finally, by substituting the equivalent values and performing a few rearrangements, we obtain:
 $\lambda^{*}=1+\frac{1}{W\left(-e^{-(\frac{k}{n\sigma^{2}}+1)}\right)}.$ (6)
 For negative real numbers in $[-1/e,0)$, Lambert function has two branches, denoted by $W_{0}$ and $W_{-1}$.
-So we need to determine that in (6) which of them should be appeared.
+Therefore, we need to determine which terms should appear in (6).
 The branch $W_{0}$ gives values between $-1$ and 00, while $W_{-1}$ gives values in $(-\infty,-1]$.
-On the other hand, due to the problem constraint $10\leq\lambda<$ 1.
-Hence, in (6) only the usage of $W_{-1}$ is valid.
-Furthermore, note that variation domain of argument of $W$ in (6) lies exactly in $[-1/e,0)$ which is domain of $W_{-1}$.
+On the other hand, due to the problem constraint $10\leq\lambda<$ 1, only the branch of $W_{-1}$ is valid in (6).
+Furthermore, note that the variation domain of the argument of $W$ in (6) lies exactly in $[-1/e,0)$, which is the domain of $W_{-1}$.
 Finally, substituting $\lambda^{*}$ in $f(\lambda)$, gives (5).
 ∎
-Despite its complicated formulation in (5), $g(c)$ has a benign manner in $c$. Figure 2 depicts that for $c>10$ its growth is highly near the linear.
+Despite its complicated formulation in (5), $g(c)$ behaves benignly with respect to $c$. Figure 2 shows that for $c>10$, its growth closely follows a linear trend.
 Figure 2: $g(c)$ versus $c$.
 Corollary III.1.
-Let $\alpha$, $\beta$ and $n$ to be given. Then $k$ is derived from $\sigma$ as follows:
+Let $\alpha$, $\beta$, and $n$ to be given. Then, $k$ is derived from $\sigma$ as follows:
 $k(\sigma;n,\alpha,\beta)=\min\left\{k_{1},k_{2}\right\},$ (7)
-where $k_{1}$ and $k_{2}$ are as follows:
+where $k_{1}$ and $k_{2}$ are defined as:
 $k_{1}(\sigma;n,\alpha,\beta)$ $\triangleq n\sigma^{2}g^{-1}\left(\frac{n\alpha^{2}\beta}{2}\right),$ (8)
@@
 Proof.
-Letting $\sigma$ vary from 00 to greater values, we see that in (3), before a critical value of $\sigma$, the first term is selected and after that, $\alpha$ is determined by the second term.
-When $\alpha$ is fixed, functionality of $k$ on $\sigma$ can be derived from one of these two terms.
-For each $\sigma$, the term that gives lower $k$ is the correct term.
+Letting $\sigma$ vary from 00 to larger values, we observe in (3) that, before reaching a critical value of $\sigma$, the first term dominates.
+After this critical value, $\alpha$ is determined by the second term.
+When $\alpha$ is fixed, the functional dependence of $k$ on $\sigma$ can be derived from either of these two terms.
+For each $\sigma$, the term that yields the smaller value of $k$ is the correct one.
 ∎
-In the following sections we set $\alpha=0.1$ and $\beta=0.05$, which are typical values in the literature. For this values, figure 2 shows that when $n>4\times 10^{4}$, the first term is highly near
+In the following sections, we set $\alpha=0.1$ and $\beta=0.05$, typical values in the literature. For these values, Figure 2 shows that when $n>4\times 10^{4}$, the first term closely approximates
 $\hat{k_{1}}(\sigma;n,\alpha,\beta)=w\cdot\frac{n^{2}\sigma^{2}\alpha^{2}\beta}{2}+bn\sigma^{2},$ (10)
@@
 IV Point-to-Point ADA in the Presence of a Gaussian Channel
-At the first step, to make our idea clear, we start with the case in which there is only one EP, answering the PS’s adaptive queries via a Gaussian channel.
-More exact, at the $i$’th question-answering round of this interactive procedure, PS sends $q_{i}(\cdot)$ through an error-free channel and receives $a_{i}$ as the answer from EP through a Gaussian channel with noise variance of $\sigma^{2}$.
-Conventionally, noise is considered as a limiting factor in the communication procedures.
-But in this work we opportunistically employ it as a constructive element.
-As noted in I-A we aim to exploit the intrinsic channel noise as the randomization noise samples in the Gaussian answering mechanism.
+To clarify our idea, we begin with the case where there is only one EP answering the PS’s adaptive queries via a Gaussian channel.
+Specifically, at the $i$’th question-answering round of this interactive procedure, the PS sends $q_{i}(\cdot)$ through an error-free channel and receives $a_{i}$ as the answer from the EP through a Gaussian channel with noise variance $\sigma^{2}$.
+As noted in I-A, we aim to exploit the intrinsic channel noise as the randomization noise samples in the Gaussian answering mechanism.
 In this scheme, at each round, such as round $i$, the EP computes the empirical average of the query as
 $q_{i}(S)=\frac{1}{n}\sum_{j=1}^{n}{q_{i}(x_{j})},$ (11)
-then multiplies it by $A_{\mathsf{t}}$.
+and then multiplies it by $A_{\mathsf{t}}$.
 The resulting signal is broadcast without additional coding, using analog Pulse Amplitude Modulation (PAM) with a maximum amplitude of $A_{\mathsf{t}}$.
-Using a matched filter, PS detects a noisy version of the transmitted signal.
-Without any hard decision, he considers the detected signal as $a_{i}$, i.e.
-the answer of his asked query.
-So we have:
+Using a matched filter, the PS detects a noisy version of the transmitted signal.
+Without applying any hard decision, it interprets the detected signal as $a_{i}$, the answer to its query.
+Therefore, we have:
 $a_{i}=A_{\mathsf{t}}q_{i}(S)+z_{i},$ (12)
 where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the sampled Gaussian channel noise at the receiver.
-Assuming that the channel condition remains unchanged during the whole ADA procedure, it’s clear that $n_{i}$’s are independent and identically distributed.
-So, the proposed scheme completely resembles the Gaussian answering mechanism in ADA and we can apply theorem III.1 to this case.
-Actually we use this theorem to set the tunable communication parameters in a way that the desired accuracy level is met.
-We assume that $(\alpha,\beta)$ has been given and the dataset size, $n$, is fixed.
-Based on this parameters and as a function of $A_{\mathsf{t}}$, we can derive the maximum number of allowable queries, for which still (2) remains valid.
-In the basic model of ADA, it’s assumed that the queries maximum value is fixed and equal to 1.
-There, the variance of the artificial randomization noise is calibrated in a way that the best performance is achieved.
-On the other hand, in our proposed model, the noise variance is fixed and determined by the physical later conditions.
-Hence, here we instead tune the queries maximum amplitude, i.e.
-$A_{\mathsf{t}}$, for the best performance achievement.
-Also note that $\alpha$ in (2) is considered under the assumption that $|q_{i}(.)|\leq 1$.
-Nonetheless, we keep using (2) by normalizing the answers maximum values to 1, when investigating this relation.
-More exact, the validity condition of (2) for the situation $(\alpha=\alpha_{0}\cdot A_{0},\beta=\beta_{0},A_{\mathsf{t}}=A_{0},\sigma=\sigma_{0})$ is completely equivalent with the situation $(\alpha=\alpha_{0},\beta=\beta_{0},A_{\mathsf{t}}=1,\sigma=\sigma_{0}/A_{0})$.
-So, by tuning $A_{\mathsf{t}}$ we implicitly calibrate the noise variance in the equivalent normalized model.
+Assuming that the channel condition remains unchanged during the whole ADA procedure, it is clear that $n_{i}$’s are independent and identically distributed.
+Thus, the proposed scheme completely resembles the Gaussian answering mechanism in ADA, allowing us to apply Theorem III.1 to this scenario.
+We use this theorem to adjust the tunable communication parameters such that the desired accuracy level is achieved.
+We assume that $(\alpha,\beta)$ has been given, and the dataset size, $n$, is fixed.
+Based on these parameters and as a function of $A_{\mathsf{t}}$, we can derive the maximum number of allowable queries for which (2) remains valid.
+In the basic ADA model, the maximum value of the queries is assumed to be fixed at 1, and the variance of the artificial randomization noise is adjusted to achieve optimal performance.
+In contrast, in our proposed model, the noise variance is dictated by the physical layer conditions and is beyond our control.
+Instead, we tune the maximum amplitude of the query, denoted as $A_{\mathsf{t}}$, to achieve the best performance.
+Also, note that $\alpha$ in (2) is considered under the assumption that $|q_{i}(\cdot)|\leq 1$.
+Nonetheless, we continue using (2) by normalizing the maximum values of the answers to 1 when investigating this relation.
+More specifically, the validity condition of (2) for the situation $(\alpha=\alpha_{0}\cdot A_{0},\beta=\beta_{0},A_{\mathsf{t}}=A_{0},\sigma=\sigma_{0})$ is entirely equivalent to the situation $(\alpha=\alpha_{0},\beta=\beta_{0},A_{\mathsf{t}}=1,\sigma=\sigma_{0}/A_{0})$.
+Thus, by tuning $A_{\mathsf{t}}$, we implicitly calibrate the noise variance in the equivalent normalized model.
 Therefore, to apply the theorem III.1, it’s sufficient to replace $\sigma$ in (7), with $\frac{\sigma}{A_{\mathsf{t}}}$.
@@
 Figure 3: Achievable number of accurately answerable queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
-Also, in figure 4 the maximum achievable number of queries is drawn versus $n$.
-In this figure, for each $n$, the best $\sigma/A_{\mathsf{t}}$ that maximizes $k$ is considered.
-As depicted in the figure, $k_{m}ax$ follows a quadratic growth in $n$.
-This observation confirms the well known fact in the ADA problem that for the worst case setup, till $O(n^{2})$ number of queries can be $(\alpha,\beta)$-accurately be answered [20].
+Additionally, in Figure 4, the maximum achievable number of queries is plotted against $n$.
+For each value of $n$, the optimal $\sigma/A_{\mathsf{t}}$ that maximizes $k$ is selected.
+As shown in the figure, $k_{\text{max}}$ exhibits quadratic growth with respect to $n$.
+This observation aligns with the well-known result in the ADA literature that, in the worst-case setup, up to $\mathcal{O}(n^{2})$ queries can be answered $(\alpha,\beta)$-accurately [20].
 Figure 4: Maximum achievable number of accurately answerable queries versus $n$ for $\alpha=0.1$ and $\beta=0.05$.
 Furthermore, Figure 3 shows that as $n$ increases, the optimum value of $\sigma/A_{\mathsf{t}}$ very slowly shifts to the left.
-This optimum point is the intersection of the quadratically increasing left curve and the exponentially decreasing right curve, , see $k_{1}$ and $k_{2}$ in (8).
-For our considered dataset sizes, optimum $\sigma/A_{\mathsf{t}}$ approximately lies in $[0.008,0.01]$.
-Furthermore, when $\sigma/A_{\mathsf{t}}$ tends to zero, the number of accurate queries reduce to values close to zero, too.
-In this low-noise regime the amount of randomization is not sufficient for ensuring that the dataset will not leak.
-Consequently, in this case by asking only few queries the analyst can overfit to the dataset and violate the accuracy criterion (2).
-Hence, in contrast to ordinary communication scenarios, increasing the power transmission may cause non-efficiency in this setup.
+This optimal point corresponds to the intersection of the quadratically increasing left curve and the exponentially decreasing right curve, as illustrated by $k_{1}$ and $k_{2}$ in (8).
+For the dataset sizes considered in this study, the optimal value of $\sigma/A_{\mathsf{t}}$ approximately falls within the range $[0.008,0.01]$.
+Furthermore, as $\sigma/A_{\mathsf{t}}$ approaches zero, the number of accurate queries also declines to near-zero levels.
+In this low-noise regime, the level of randomization becomes insufficient to prevent data leakage.
+As a result, the analyst can overfit to the dataset and violate the accuracy criterion in (2) with only a few queries.
+Thus, unlike standard communication scenarios, increasing transmission power in this setup can lead to inefficiency.
 On the other hand, computation shows that in order to have a non-zero number of $(\alpha,\beta)$-accurate query-answers by this scheme, $\sigma/A_{\mathsf{t}}$ should be lower than 0.017 or equivalently $A_{\mathsf{t}}$ should be approximately 60 times greater than the standard deviation of the noise.
-This ratio for the optimum situation lies in $[100,125]$.
-Equivalently, a minimum SNR of 32 dB is required and optimum SNR is in the range [37, 38.9] dB.
-However, in high-noise regime, considering practical limitations on the transmission power, improving $k$ may not be straight-forward.
-We show in the next section that federation of some edge answering points can remedy this situation.
+This ratio, in the optimum situation, lies within the range of $[100,125]$.
+Equivalently, a minimum SNR of 32 dB is required, with the optimum SNR falling between 37 and 38.9 dB.
+However, in the high-noise regime, given practical limitations on transmission power, improving $k$ may not be straightforward.
+In the next section, we demonstrate how a federation of edge answering points can address this issue."
 V Distributed ADA in the Presence of a Gaussian Channel
-In the previous section we considered the point-to-point scenario to make the idea of usefully exploiting the channel noise clear.
-We showed that by calibrating the amplitude of transmitted signal, achievable number of accurately answered queries can be optimized.
+In the previous section, we examined the point-to-point scenario to clarify the concept of effectively exploiting the channel noise.
+We showed that calibrating the amplitude of the transmitted signal can optimize the achievable number of accurately answered queries.
 Furthermore, we see that optimum $A_{\mathsf{t}}$ depends on the dataset size, $n$.
-On the other hand, although we’ve succeeded theoretically to utilize the noise as a part of adaptive query answering procedure, but the required SNR range may not be manageable in some of communication situations.
-This concern becomes more crucial when the answerer is a personal device or a low-power sensor at the edge of network with a long distance from the receiver.
-As we mentioned previously, this can be a reasonable use case of our proposed setup.
-In addition, we observed that for a fixed $A_{\mathsf{t}}$ the required dataset size, i.e.
-$n$, grows by $O(\sqrt{k})$.
-So, to guarantee more accurate (or private) rounds of data analysis, the edge point needs to answer the queries based on larger datasets.
-However, it may be impractical for him to gather or store large samples.
-In this section we show that both of the power and data limitations can be handled in the distributed scenario.
-We consider a situation that $L$ Edge Points simultaneously transmit their answers to adaptive queries of the Central point through an analog Gaussian multiple access channel (MAC).
-For simplicity, let’s assume that the dataset sizes of EPs are the same, i.e.
-$n_{l}=n_{0}$ for any $l\in\{1,2,\dots,L\}$.
+On the other hand, although we have theoretically succeeded in utilizing noise as part of the adaptive query-answering procedure, the required SNR range may not always be feasible in certain communication scenarios.
+This challenge is particularly significant when the answerer is a personal device or a low-power sensor located at the network’s edge and operating at a considerable distance from the receiver.
+As previously mentioned, this represents a plausible use case for our proposed setup.
+Additionally, we observed that for a fixed $A_{\mathsf{t}}$, the required dataset size $n$ grows at a rate of $\mathcal{O}(\sqrt{k})$.
+Consequently, ensuring more accurate (or private) rounds of data analysis would require the EP to answer queries based on increasingly larger datasets.
+However, it may be impractical for the EP to gather or store such large datasets.
+In this section, we demonstrate that both power and data limitations can be addressed within a distributed framework.
+We consider a situation where $L$ EPs simultaneously transmit their answers to the CP’s adaptive queries through an analog Gaussian multiple access channel (MAC).
+For simplicity, we assume that the dataset sizes of all EPs are identical, i.e., $n_{l}=n_{0}$ for all $l\in{1,2,\dots,L}$.
 Similar to the previous section, each EP, like $\mathcal{M}_{l}$, computes $q_{i}(S_{l})=\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}$ and transmit $A_{\mathsf{t}}q_{i}(S_{l})$ without any coding.
 Because of the additive nature of the air interface, the receiver gets a noisy version of the aggregated signals.
-So, the received signal at the CP will be $a_{i}=A_{\mathsf{t}}\sum_{l=1}^{L}{q_{i}(S_{l})}+z_{i}$, in which $z_{i}\sim\mathcal{N}(0,\sigma^{2})$.
-In order to make an equivalence with the point-to-point case, we consider $a_{i}/L$ in the following.
-By substituting $q_{i}(S_{l})$, we will have:
+So, the received signal at the CP will be $a_{i}=A_{\mathsf{t}}\sum_{l=1}^{L}{q_{i}(S_{l})}+z_{i}$, where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$.
+To establish equivalence with the point-to-point case, we consider $a_{i}/L$ in the following.
+Substituting $q_{i}(S_{l})$ yields:
 $\frac{a_{i}}{L}=\frac{A_{\mathsf{t}}}{Ln_{0}}\sum_{l=1}^{L}{\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}}+\frac{z_{i}}{L}.$ (13)
 The above equation completely resembles a point-to-point scenario with $n_{\text{eq}}=Ln_{0}$, $\sigma_{\text{eq}}=\frac{\sigma}{L}$ and $A_{T,eq}=A_{\mathsf{t}}$.
-We see that by this cooperative transmission scheme, both the equivalent dataset and the noise variance is improved by a factor of $L$.
-By considering these equivalent parameters, the previous section results can be applied in this scenario, too.
-Therefore, we can use (7) with the aforesaid $n_{\text{eq}}$ and $\sigma_{\text{eq}}$.
-In this case, we fix $\sigma/A_{\mathsf{t}}=0.5$ and derive $k$ as a function of $L$, for different values of $n$.
-Considering the previous section, we know that $L$ should be large enough to make $\sigma_{\text{eq}}$ less than $0.017$ and consequently, attain a positive value for $k$.
-Figure 5 shows the results for this scenario.
-The parameters $\alpha$ and $\beta$ are the same with the previous scenario.
+We observe that with this cooperative transmission scheme, both the equivalent dataset size and the noise variance are improved by a factor of $L$.
+Using these equivalent parameters, the results from the previous section can be applied to this scenario as well.
+Thus, we can use (7) with the corresponding $n_{\text{eq}}$ and $\sigma_{\text{eq}}$.
+In this case, we fix $\sigma/A_{\mathsf{t}}=0.5$ and derive $k$ as a function of $L$ for different values of $n$.
+Based on the results from the previous section, we know that $L$ must be sufficiently large to ensure that $\sigma_{\text{eq}}$ is less than $0.017$, thereby yielding a positive value for $k$.
+Figure 5 illustrates the results for this scenario.
+The parameters $\alpha$ and $\beta$ remain the same as in the previous case.
 Figure 5: Maximum number of accurately answered queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
-We see that for each $n$, before a critical number of EPs, $k$ grows exponentially in $L$, which is due to under-leakage regime and is related to $k_{2}$ in (8).
-This critical $L$ shifts very slowly in $n$ to the right.
-For each $n$, after this $L$, growth of $k$ becomes very slow such that in figure 5 it seems constant.
-The reason is that after the critical $L$, $k$ obeys from $k_{1}$ function in (8).
-As we noted previously, $k_{1}$ can be approximated by $\hat{k_{1}}$ in (10).
-In this function, the quadratic term is strongly dominant for large $n$.
+We observe that for each $n$, before reaching a critical number of EPs, $k$ grows exponentially with $L$.
+This is due to the under-leakage regime and is related to $k_{2}$ in (8).
+The critical value of $L$ shifts slowly to the right as $n$ increases.
+After this critical $L$, the growth of $k$ becomes very slow, and in Figure 5, it appears nearly constant.
+This is because, beyond the critical $L$, $k$ follows the $k_{1}$ function from (8).
+As mentioned earlier, $k_{1}$ can be approximated by $\hat{k_{1}}$ in (10).
+In this function, the quadratic term becomes strongly dominant for large $n$.
 In this quadratic term, $n^{2}\sigma^{2}$ is determining.
-On the other hand, in the distributed scenario we have
-$n_{\text{eq}}\sigma_{\text{eq}}=(Ln)(\sigma/L)=\sigma,$
+On the other hand, in the distributed scenario, we havebb
+$n_{\text{eq}}\sigma_{\text{eq}}=(Ln)(\sigma/L)=\sigma,$ (14)
 which does not depend on $L$. This argument also provides a perspective for addressing this inefficient growth stoppage, which is discussed in the following subsection.
-V-A Improving number of queries in EP counts
-Considering our descriptions on the figure 3 about the optimum value of $\sigma/A_{\mathsf{t}}$, EPs in the distributed scenario need to decrease $A_{\mathsf{t}}$ with respect to $L$, after that $L$ passes the critical value.
-For large $n_{\text{eq}}$, optimum $\sigma_{\text{eq}}/A_{\mathsf{t}}$ is about $0.008$.
+V-A Improving $k$ in $L$
+Considering the descriptions in Figure 3 about the optimum value of $\sigma/A_{\mathsf{t}}$, EPs in the distributed scenario need to decrease $A_{\mathsf{t}}$ as $L$ exceeds the critical value.
+For large $n_{\text{eq}}$, the optimum value of $\sigma_{\text{eq}}/A_{\mathsf{t}}$ is approximately $0.008$.
 The exact value can be derived by setting $k_{1}=k_{2}$ in (7).
-Naming this value by $s_{\text{opt}}(n_{\text{eq}})$, optimum $\sigma/A_{\mathsf{t}}$ will be $Ls_{\text{opt}}(n_{\text{eq}})$.
-In other words, EPs should tune their $A_{\mathsf{t}}$ equal to $\frac{\sigma}{Ls_{\text{opt}}(Ln)}$.
-As we noted, in the previous section, $s_{\text{opt}}$ changes very slowly.
-Hence, we can approximately say that in optimum regime $A_{\mathsf{t}}\propto 1/L$.
-Figure 6 depicts the resulted improvement in the growth of $k$, after optimizing $A_{\mathsf{t}}$.
+Denoting this value as $s_{\text{opt}}(n_{\text{eq}})$, the optimum value of $\sigma/A_{\mathsf{t}}$ will be $Ls_{\text{opt}}(n_{\text{eq}})$.
+In other words, EPs should tune their $A_{\mathsf{t}}$ to $\frac{\sigma}{Ls_{\text{opt}}(Ln)}$.
+As noted in the previous section, $s_{\text{opt}}$ changes very slowly.
+Therefore, we can approximately state that in the optimum regime, $A_{\mathsf{t}}\propto 1/L$.
+Figure 6 illustrates the resulting improvement in the growth of $k$ after optimizing $A_{\mathsf{t}}$.
 Figure 6: Maximum number of accurately answered queries versus $\sigma/A_{\mathsf{t}}$ for $\alpha=0.1$ and $\beta=0.05$ and different values of $n$.
 VI Future Work and Conclusion
-In this work, after making a previous result in ADA more proper for our analyses, we showed that Gaussian channel noise can be used as an opportunity when the basic model of ADA is considered.
-Of course, in this situation minimum required SNR was considerable.
-Then, we showed that the federation of the edge answering points can improve the equivalent noise variance and total dataset size by a linear factor when the shared medium is a Gaussian MAC.
-One important extension for further study can be considering the fading channel, which is a more realistic model.
-In this situation, each EP’s transmission experiences a multiplicative distortion before receiving by the CP.
-In the other word, the received signal during the i’th answering round will be as follows
+In this work, after refining a previous result in ADA to better suit our analyses, we showed that Gaussian channel noise can be leveraged as an opportunity when the basic model of ADA is considered.
+Of course, in this situation, the minimum required SNR was significant.
+Then, we demonstrated that the federation of edge answering points can linearly improve the equivalent noise variance and total dataset size when the shared medium is a Gaussian MAC.
+One important extension for further study could involve considering the fading channel, which represents a more realistic model.
+In this case, each EP’s transmission undergoes a multiplicative distortion before being received by the CP.
+In other words, the received signal during the $i$-th answering round would be as follows:
 $Y_{i}=\sum_{l=1}^{L}{h_{il}X_{il}+z_{i}},$ (15)
-in which $X_{il}$ is the transmitted signal by the $l$’th EP and $h_{il}$ is the fading coefficient of the wireless channel from this EP to the CP.
-Based on various scenarios of Channel State Information (CSI) knowledge, different scenarios can be studied: CSI known only at the transmitter, CSI known only at the receiver, and CSI not known at either.
-Overally, a main challenge in the fading scenario is that the unweighted aggregation of transmissions is not intrinsically prepared anymore.
+in which $X_{il}$ represents the signal transmitted by the $l$-th EP, and $h_{il}$ denotes the fading coefficient of the wireless channel between this EP and the CP.
+Based on the availability of CSI, various scenarios can be analyzed: CSI known only at the transmitter, CSI known only at the receiver, and CSI unknown at both ends.
+A primary challenge in the fading scenario is that the unweighted aggregation of transmissions is no longer inherently feasible.